Suppose that an annuity will provide for 20 annual payments of 1280 dollars,with the first payment coming 10 years from now.if the nominal rate of interest is 10.6 percent convertible monthly,what is the present value of the annuity?


1
Expert's answer
2021-05-07T11:45:32-0400

P=PMT*( 1−(1/(1+r)^n))/r

where:

P=Present value of an annuity stream

PMT=Dollar amount of each annuity payment

r=Interest rate (also known as discount rate)

n=Number of periods in which payments will be made

P=(1280*20) *(1 -(1/1.106^10)/0.106

= 25,600 * 5.98933

= 153,326.848